A large tank is filled to capacity with 700 gallons of pure water. Brine containing 2 pounds of salt per gallon is pumped into the tank at a rate of 7 gal/min. The well-mixed solution is pumped out at a rate of 14 gals/min. Find the number A(t) of pounds of salt in the tank at time t.



1
Expert's answer
2021-10-20T01:21:35-0400

"\displaystyle
\frac{dA}{dt} = R_{in}-R_{out}\\
R_{in} = (2lb/gal \cdot 7gal/min) =14 lb/min\\
R_{out} = (\frac{A(t)}{700}lb/gal).(14gal/min)=\frac{A(t)}{50}lb/min\\
\text{Hence, } \frac{dA}{dt} = 14- \frac{A(t)}{50}\\
\text{Integrating, we have that}\\
\ln |14- \frac{A}{50}| = -0.02t +\ln c\\
14-\frac{A}{50} = ce^{-0.02t}\\
A(t) = 700 - 50ce^{-0.02t}\\
\text{Applying the initial conditions, we have}\\
A(0) = 700 - 50ce^0=0\\
\implies c = 50\\
\therefore A(t) = 700 - 700e^{-0.02t}"